OpenStudy (anonymous):

what are three equivalent fractions of 4/9?

OpenStudy (anonymous):

Just multiply the numerator and denominator by 1. But not just any 1, like 2/2, 3/3, 4/4, etc. Why? bc they all equal to 1 once you reduce them. So, 4/9 times 2/2 = 8/18 4/9 times 3/3 = 12/27 4/9 times 4/4 = 16/36 etc...

OpenStudy (anonymous):

thank you veronometry! I'm trying to help my daughter but i have forgotten all I learned